A PRIMER ON DUNHAM’S FORMULATION OF VIBRATION-ROTATIONAL ENERGIES OF DIATOMIC MOLECULES

\underline{J. F. OGILVIE}, Universidad de Costa Rica, Escuela de Quimica, Ciudad Universitaria Rodrigo Facio, San Pedro de Montes de Oca, San Jose 2060, Costa Rica; also, Simon Fraser University, Department of Mathematics and Centre for Experimental and Constructive Mathematics, 8888 University Drive, Burnaby, British Columbia V5A 1S6, Canada.

The formulation by J. L. Dunham (1932) of a systematic approach to representation of the vibration-rotational energies of diatomic molecules in discrete states, and of intensities of transitions between such states, is much better known by name than understood. Pertinent aspects, including the primacy of Dunham coefficients, are reviewed.

J. L. Dunham, Physical Review, 41, 713 -720, 721-731 (1932)